OpenStudy (potatobibin):

You are given a number of 10 Ω resistors, each capable of dissipating only 3.0 W without being destroyed. What is the minimum number of such resistors that you need to combine in series or in parallel to make a 10 Ω resistance that is capable of dissipating at least 8.6 W? My answer is 1 which is wrong idk why pls help. Tried searching it.. I got wrong

OpenStudy (mrnood):

Power = I^2R so what current is required, through 10 ohms to give 8.6W?

OpenStudy (potatobibin):

Okay then?

OpenStudy (potatobibin):

The question is kinda tricky for me. So maybe I got wrong in the reading comprehension what I did is 8.6=(x/10)^(-1) and I got wrong

OpenStudy (mrnood):

I'll rethink this ... The resultant resistance is required to be 10 ohms you start with 10 ohm resistors if you put them in parallel you reduce the resistance if you put them in series you increase the resisistance so - can you deduce the layout of the combination that will remain 10 ohms?

OpenStudy (potatobibin):

Hmm I getting lost sir.. can you please guide me more.

OpenStudy (mrnood):

you are require to make a combination of reistors the result of the combination must be 10 ohms but each individual resitor is 10 ohm so for instance if you put 2 in parallel you reduce the resistance to 5ohms if you put 2in series you increase the resisistance to 20 ohms so how would you arrange the resistors so that the resultant resistance is 10 ohms?

OpenStudy (potatobibin):

So you only need 1?

OpenStudy (potatobibin):

Uh wait hmm

OpenStudy (potatobibin):

2 parallel resistors and 0 series resistors?

OpenStudy (mrnood):

no - because 1 can only withastand 3W power you need to make a combination that can dissipate at least 8.6W BUT the combination must also have a resistance of 10 ohms

OpenStudy (mrnood):

if you put 2 in parallel you reduce the resistance to 5ohms

OpenStudy (potatobibin):

I need to get a 9?

OpenStudy (potatobibin):

it dissipates 8.6 and less than 10

OpenStudy (potatobibin):

or wait I need 3 resistors?

OpenStudy (mrnood):

do you understand the rlationship between current ,resistance and power? also the parallel and series resistance equations?

OpenStudy (potatobibin):

I know the formula... which is kinda a problem maybe? I'm so used on calculating that I forgeting the true concept of the topic

OpenStudy (mrnood):

so - your guess of 3 cannot be true - how would you make 10 ohms resultant from 3 10 ohm resistors?

OpenStudy (potatobibin):

oh my I'm so getting lost..

OpenStudy (mrnood):

so using W = I^2 R you need to work out th current require dto get 8.6 W through 10 ohms AND The maximaum current that can pass through 1 of the original resistors

OpenStudy (potatobibin):

Okay so by that formula I can get it? HMm wait

OpenStudy (mrnood):

No - that is the starting point you need to know the current before you can work out the answer

OpenStudy (potatobibin):

W = 8.6 right

OpenStudy (potatobibin):

then R = 10?

OpenStudy (mrnood):

yes = so work out I from W = I^2 R

OpenStudy (potatobibin):

So I got I = 0.9274 what's next?

OpenStudy (mrnood):

OK - so that is the minimum current that must pass through the combination of resisstors, in order to get at least 8.6W So NOW What is the maximum current that can pass through 1 of the original 10 ohm resistors if the power is 3W

OpenStudy (mrnood):

so work out I from W = I^2 R

OpenStudy (potatobibin):

Okay maximum eh.. Hmmm thinking

OpenStudy (mrnood):

maximum current is when power is maximum - so max current is when w=3, r=10

OpenStudy (potatobibin):

will I change 8.6W into 3W?

OpenStudy (potatobibin):

3=(I)^2(10) ??

OpenStudy (mrnood):

yes

OpenStudy (potatobibin):

I goy 0.55

OpenStudy (mrnood):

ok - correct SO you can see now why you cannot use 1 resistor - the current is too high - it will burn out so  if the 2 resistances are equal then how much current flows through each resisttance?